OpenStudy (anonymous):

if sally can paint a room in 4 hours alon and stever can paint a room in 5 hours alone. how long will it take for them to paint a room together?

OpenStudy (watchmath):

Sally speed =1/4 (means 1 room in 4 hours). Steve speed = 1/5. Together the speed become 1/4 + 1/5 = 9/20. So 9 room in 20 hours. For a room they just need 20/9 hours.

OpenStudy (anonymous):

thank you!!!!!!